# You Only Learn One Query: Learning Unified Human Query for Single-Stage Multi-Person Multi-Task Human-Centric Perception

### 09 Dec 2023 v1

@@ 4.2 Task-shared Transformer Decoder
Model Backbone Det. Seg. Kpt. Cls. (Gender) Cls. (Age)
@@
$\bigstar$ HQNet (Ours) R-50 74.9 70.4 80.7 65.8 58.7 73.9 69.3 63.8 77.3 56.0 42.5 63.3 53.8 39.7 61.2
$\bigstar$ HQNet (Ours) Swin-L 77.3 73.3 82.7 68.1 60.9 75.9 72.6 67.4 80.1 57.9 43.1 65.8 56.2 41.5 63.9
$\bigstar$ HQNet (Ours) ViT-L 78.0 73.6 83.7 68.6 61.4 76.5 75.3 69.8 83.5 58.0 44.7 65.0 58.0 40.9 66.7
4.3 Task-specific Heads
To ensure scalability, we categorize HCP tasks into three groups and design specific implementation paradigms for each category. Coordinate prediction tasks (e.g. object detection and pose estimation) share common reference points with bounding box prediction and directly regress the normalized offsets of each point. Dense prediction tasks (e.g. instance segmentation and human parsing) follow the design of Mask DINO [41], which involves constructing a high-resolution pixel embedding map by integrating features from both the backbone and the Transformer encoder. By performing a dot-product operation between the content query embedding and the pixel embedding map, an instance-aware pixel embedding map is generated, facilitating pixel-level classification. Classification tasks (e.g. determining if an instance is human, gender and age estimation) directly map the Human Query to the classification prediction results, as the Human Query inherently encodes the positional information.
To minimize the overhead of incorporating new tasks, we employ lightweight task-specific heads comprising simple linear layers. The details of our four task-specific heads are as follows: Human detection head. A 3-layer multi-layer perceptron (MLP) with a hidden dimension of $d$ is utilized to predict the normalized center x-y coordinates, height, and width of the bounding box w.r.t. the input image. Additionally, a linear projection layer (FC) is employed to predict the class label (human or non-human). Pose estimation head. Following the coordinate prediction paradigm, the learned Human Query is fed into a pose regression head (MLP) to regress the relative pose offsets w.r.t. the shared reference points of the detection head. A confidence prediction head (FC) is used to predict confidence score of having visible keypoints. Following PETR [72], joint decoder layers are employed to refine body poses by leveraging structured relations between body keypoints. An auxiliary heatmap branch is used to aid training and discarded during testing. Human instance segmentation head. A 3-layer MLP is used to process the instance-aware pixel embedding map and output a one-channel mask, which is then upsampled to match the original input image size. Human attribute head. The gender estimation head and the age estimation head operate in parallel. Both heads consist of two-layer MLPs. Gender estimation involves binary classification, while age estimation is formulated as an 85-class ([1, 85]) classification with softmax expected value [68] estimation.
5 Experiments
@@ 5.1 Dataset and Evaluation Metric
COCO-UniHuman Dataset. Our model training exclusively employs COCO-UniHuman train data (in addition to ImageNet pre-training). We follow DINO [96] for augmentation and adopt the 100-epoch training schedule. Model evaluation takes place on COCO-UniHuman val set (5K images). Due to limitations of the COCO test-dev evaluation server, which lacks support for “Person” category evaluation and attribute recognition, we mainly report results on the val set. The evaluation is based on the standard COCO metrics including Average Precision (AP), $\text{AP}^{M}$ for medium-sized persons and $\text{AP}^{L}$ for large-sized persons. Following [100, 101], we exclude the Small category persons during evaluation due to the lack of annotations in COCO. For attribute recognition, we also use AP with Age-10 metric for evaluation, where the age estimation is considered correct if the prediction error is no larger than 10.
OCHuman Dataset. OCHuman dataset is a large benchmark dataset that focuses on heavily occluded humans. It contains no training samples and is intended solely for evaluation purposes. Following [100], we train models on the COCO train set and evaluate models on OCHuman val set (4731 images) and test set (8110 images).
5.2 Results on COCO dataset
We compare our method to task-specific and multi-task HCP models on the COCO-UniHuman dataset in Table 2. It outperforms multi-task HCP models and achieves very competitive results against task-specific HCP models. More details about the baselines can be found in Sec. A5.
@@ 5.2 Results on COCO dataset
Det. Seg. Kpt. Det. Seg. Kpt.
$\diamondsuit$ Mask R-CNN R-50-FPN - 16.3 - 16.9
$\diamondsuit$ ${}^{\dagger}$ R-50 37.8 30.4
$\diamondsuit$ Pose2Seg R-50-FPN 22.2 28.5 23.8 30.3
$\heartsuit$ ${}^{\dagger}$ HRNet-w32 37.9 36.5
$\heartsuit$ ${}^{\dagger}$ HRNet-w32 40.0 39.4
$\bigstar$ YOLACT R-101-FPN 13.2 13.5
$\bigstar$ CondInst R-50-FPN 20.3 20.1
$\bigstar$ LSNet-5 DLA-34 25.0 24.9
$\bigstar$ ${}^{\dagger}$ HRNet-w32 39.0 38.1
$\bigstar$ ${}^{\dagger}$ R-50-FPN 29.2 28.3
$\bigstar$ ${}^{\dagger}$ HRNet-w32 44.9 44.0
@@
Comparison with multi-task HCP methods. Pose2Seg [100] is a two-stage human pose-based instance segmentation approach. It uses a standalone keypoint detector for pose estimation and employs human skeleton features for top-down instance segmentation guidance. MultiPoseNet [1] and PersonLab [65] follow bottom-up strategies. CenterNet [105], LSNet [101], and UniHead [45] [^3] UniHead trains separate models for different HCP tasks. are single-stage alternatives. Our R-50 model achieves superior detection, keypoint estimation, and segmentation performance in multi-task HCP, without bells and whistles.
@@ 5.3 Results on the OCHuman dataset
To verify the performance of HQNet in more challenging crowded scenarios, we compare it with recent works on the OCHuman benchmark, which is a popular crowded scene benchmark for human detection, segmentation, and pose estimation. We show that our model outperforms previous methods under the same ResNet50 backbone network by a large margin. For instance, it outperforms SBL by 9.6 keypoint AP and CondInst by 11.0 segmentation AP on test set. It even achieves superior performance than HrHRNet (40.3 vs 40.0) and LOGO-CAP (40.3 vs 39.0) even with a much smaller backbone (ResNet-50 vs. HRNet-w32). With a stronger backbone, i.e. ViT-L, our HQNet sets new state-of-the-art results on detection (35.8 AP), segmentation (38.8 AP), and pose estimation (45.6 AP).
@@
Table 5: Unseen-task (MOT) evaluation on PoseTrack21 [15]. ‘FT’ means fine-tuning on PoseTrack21. Our models are evaluated without training on MOT.
@@
Table 5: Unseen-task (MOT) evaluation on PoseTrack21 [15]. ‘FT’ means fine-tuning on PoseTrack21. Our models are evaluated without training on MOT.
@@
Table 7: Analysis of multi-task co-learning on COCO-UniHuman val set. We report AP for the “Person” category without Small category person. All models use ResNet-50 backbone with 100-epoch training setting for fair comparisons.
Model Params Det. Seg. Kpt. Cls. (Gender) Cls. (Age)
Det. 47.29 M 73.3 68.1 79.8
Det.+Seg. 48.30 M 73.0 68.0 79.4 63.6 57.6 72.1
Det.+Kpt.+Seg. 51.99 M 74.5 70.3 80.1 65.7 58.7 73.8 69.5 64.4 77.0
Det.+Kpt.+Seg.+Cls. 52.14 M 74.9 70.4 80.7 65.8 58.7 73.9 69.3 63.8 77.3 53.8 39.7 61.2 56.0 42.5 63.3
@@ 5.6 More Analysis
Multi-task co-learning. In Table 7, we observed that co-learning with multiple human-centric tasks leads to improved overall performance. This enhancement can be attributed to the inter-task synergy that arises from jointly training different HCP tasks.
Computational cost analysis. In Table 7, we follow [72] to analyze the computational cost. Specifically, we use ResNet-50 backbone with the input size of $800\times 800$. In HQNet, multiple tasks share the computation cost of the backbone, transformer encoder and decoder. The overhead of each task-specific head is negligible, showing good scalability of HQNet in terms of increasing the number of tasks.
Effect of HQ-Ins Matching. As shown in Figure 3 (left), with box only matching [96], there may be some erroneous cases when one person’s pose is matched to another person. HQ-Ins Matching avoids such errors by comprehensively considering multiple tasks as a whole. Quantitative evaluation can be found in Sec. A3.2.
Figure 3: Effect of HumanQuery-Instance (HQ-Ins) Matching. w/o HQ-Ins Matching: Using detection loss only for bipartite matching [96]. w/ HQ-Ins Matching: Comprehensively using detection, keypoint, and segmentation loss for bipartite matching.
@@ 6 Conclusion
Limitations. While we focus on 2D tasks, tasks about 3D [58, 59, 86] or sequential data [21, 89] also hold significant potential. We encourage future research to explore more comprehensive multi-task human-centric perception.
Acknowledgement. We thank Wang Zeng and Ruijie Yao for valuable discussions. This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
@@
Figure A1: Statistics of COCO-UniHuman benchmark. (a) The gender distribution of COCO-UniHuman is biased towards male. (b) The age distribution ranges from [1, 84] and is biased towards young adults, since images are from public Internet repositories.
In Figure A1, we show statistics of our proposed COCO-UniHuman dataset. The plots show the distribution of the gender and the apparent age. We find gender and age biases existed in the widely used COCO dataset. The occurrence of men is significantly higher than women in COCO dataset. More specifically, male to female ratio is about 65:35. In addition, the dataset has an unbalanced age distribution. The apparent age distribution ranges from [1, 84], and it is mainly concentrated between the ages of 25 and 35. Analyzing and addressing the gender and age bias in the computer vision system can also be an important topic in the AI community. Future work could also use our benchmark dataset to comprehensively measure and analyze such biases, but it is out of the scope of this paper.
@@
Multi-task co-learning can mitigate over-fitting. From Figure A2, we observe that training task-specific models on “Person” category only will easily lead to over-fitting problem, the performance decreases with increasing number of epochs. Specifically, in this experiment, we compare the common 1x, 2x, and 4x training settings for RCNN-based methods (i.e. Faster-RCNN, Mask RCNN), and compare 50-epoch and 100-epoch settings for DETR-based methods (i.e. DINO, Mask DINO, and our HQNet). The models are trained using MMDetection [9] with suggested hyper-parameters. We report the Average Precision (AP) for both human detection (solid lines) and the human instance segmentation (dashed lines) on COCO-UniHuman val set. Interestingly, we find that our presented multi-task co-learning (HQNet) can mitigate the over-fitting problem, and the performance consistently improves with the increasing training epochs, demonstrating good scalability.
@@
A3.1 General models vs person-specific models
@@
Table A4: Effect of HumanQuery-Instance (HQ-Ins) Matching. Experiments are conducted on COCO-UniHuman val set using the ResNet-50 backbone with 100-epoch training setting. We report AP for the “Person” category without Small category person.
Matching Det. Seg. Kpt. Cls. (Gender) Cls. (Age)
@@
In Figure A3 and Figure A4, we show some qualitative results of HQNet with ResNet-50 backbone. In Figure A3, we show some qualitative results on COCO-UniHuman val dataset for human detection, human pose estimation, human instance segmentation, and human attribute recognition. The model can recognize the gender and age of different people
In Figure A4, we visualize the results of human detection and tracking (same color for same id), human pose estimation, human instance segmentation, gender estimation and age estimation. As introduced in the section of “Unseen-task generalization” in the main paper, we directly apply our HQNet on multiple object tracking (with pose estimation and segmentation) on the challenging PoseTrack21 [15] dataset, where our models are trained only on the COCO-UniHuman image-based dataset without explicitly tuned for multi-object tracking (MOT) on video-based dataset like PoseTrack21. Note that our results are simply obtained by image-based human analysis without performing further smoothing post-process. Our learned Human Query, which encodes both spatial and visual cues, can serve as good embedding features to distinguish different human instances. Therefore, our human tracking is robust to heavy occlusion, and the id can recover from occlusions. Our HQNet makes a comprehensive all-in-one human analysis system that can achieve multiple functions: multiple object tracking with human pose estimation, human instance segmentation, and human attribute recognition.
@@
In Figure A5, we visualize the sampling locations of deformable attention for different HCP models. We show the results of the last decoder layer in HQNet-ResNet50. Each sampling point is marked as a red-filled circle. The left results are from the model trained for detection and segmentation ($M_{Det.+Seg.}$). The middle ones are from the model trained for detection and keypoint ($M_{Det.+Kpt.}$). And the right ones are from the model trained for detection, segmentation, keypoint and attribute ($M_{Det.+Kpt.+Seg.+Cls.}$). With the segmentation task, we notice that some of the sampling points of $M_{Det.+Seg.}$ are distributed near the boundary of the human body, and some are distributed in the background to capture more context information. The sampling points of $M_{Det.+Kpt.}$ have higher probability to distribute inside the human body and some of the points are located closer to the defined human body keypoints, especially the face, arms, and legs. $M_{Det.+Kpt.+Seg.+Cls.}$ combines the characteristics of $M_{Det.+Seg.}$ and $M_{Det.+Kpt.}$.
Figure A3: Qualitative results on COCO-UniHuman val dataset. Our HQNet achieves accurate human detection, human pose estimation, human instance segmentation, and human attribute recognition simultaneously.
Figure A4: Qualitative results on PoseTrack21 val set. Our HQNet makes a comprehensive human analysis system that can achieve multiple functions: multiple object tracking with pose estimation, human instance segmentation, and attribute recognition. Note that our HQNet is only trained on the COCO-UniHuman image-based dataset without finetuning on the PoseTrack21 video-based dataset.
Figure A5: Visualization of deformable attention sampling points. The results are from models trained for different HCP tasks. Left: detection and segmentation. Middle: detection and keypoint. Right: detection, keypoint, segmentation and attribute.
@@
In this work, we jointly train multiple human-centric perception (HCP) tasks, including human detection, human instance segmentation, human pose estimation and human attribute (gender and age) recognition.
For human detection, we follow DINO [96] to apply focal loss [48] for classification $L^{focal}_{cls}$ and detection loss (L1 regression loss $L^{reg}_{det}$ and GIOU loss [67] $L^{giou}_{det}$). For human pose estimation, we follow PETR [72] to use focal loss for classifying valid and invalid human instances $L^{focal}_{kpt}$, L1 keypoint regression loss $L^{reg}_{kpt}$, OKS loss $L^{oks}_{kpt}$, and auxiliary heatmap loss $L^{hm}_{kpt}$. For segmentation loss, we use binary cross-entropy loss $L^{bce}_{seg}$ and dice loss $L^{dice}_{seg}$. For attribute recognition, we have binary cross-entropy loss for gender estimation $L^{bce}_{gender}$ and mean-variance loss [64] for age estimation $L^{mean}_{age}$ and $L^{var}_{age}$.
@@
$+\lambda^{focal}_{kpt}L^{focal}_{kpt}+\lambda^{reg}_{kpt}L^{reg}_{kpt}+\lambda^{oks}_{kpt}L^{oks}_{kpt}+\lambda^{hm}_{kpt}L^{hm}_{kpt}$
@@
$+\lambda^{bce}_{gender}L^{bce}_{gender}+\lambda^{mean}_{age}L^{mean}_{age}+\lambda^{var}_{age}L^{var}_{age},$
@@
Keypoint $\lambda^{focal}_{kpt}$ 1.0
@@
We follow the setting of DINO [96] to augment the input image by random crop, random flip, and random resize. Specifically, we randomly resize the input image to have its shorter side between 480 and 800 pixels and its longer side less or equal to 1333.
The models are trained with AdamW optimizer [35] with base learning rate of $1\times 10^{-4}$, momentum of $0.9$ and weight decay of $1\times 10^{-4}$. For all experiments, the models are trained for 100 epochs with a total batch size of 16 and the initial learning rate is decayed at 80th epoch by a factor of 0.1. We use 16 Tesla V100 GPUs for model training.
@@
Details about human instance segmentation baselines. For human instance segmentation, we contrast HumanQuery with state-of-the-art general and human-specific instance segmentation methods. Mask R-CNN [24] is an end-to-end top-down approach that optimizes object detection and instance segmentation jointly. Given our one-stage pipeline, we also compare against one-stage methods , including PolarMask [88], MEInst [98], YOLACT [5], and CondInst [77]. Results of PolarMask, MEInst, YOLACT, CondInst are from [101], which are obtained by re-training and evaluating the models on COCO “Person” category only. PolarMask encodes the instance mask with coordinates, while MEInst encodes the mask into a compact representation vector. YOLACT and CondInst use a series of global prototypes and linear coefficients to represent instance masks. Instead we learn instance-aware Human Query to decouple each human instance.

### 14 Mar 2024 v2

@@ 5 Experiments
Model Backbone Det. Seg. Pose (Kpt.) Cls. (Gender) Cls. (Age)
@@
$\bigstar$ HQNet (D+S) R-50 73.0 68.0 79.4 63.6 57.6 72.1
$\bigstar$ HQNet (D+S+P) R-50 74.5 70.3 80.1 65.7 58.7 73.8 69.5 64.4 77.0
$\bigstar$ HQNet (D+S+P+C) R-50 74.9 70.4 80.7 65.8 58.7 73.9 69.3 63.8 77.3 56.0 42.5 63.3 53.8 39.7 61.2
$\bigstar$ HQNet Swin-L 77.3 73.3 82.7 68.1 60.9 75.9 72.6 67.4 80.1 57.9 43.1 65.8 56.2 41.5 63.9
$\bigstar$ HQNet ViT-L 78.0 73.6 83.7 68.6 61.4 76.5 75.3 69.8 83.5 58.0 44.7 65.0 58.0 40.9 66.7
Table 3: Results of 3D human mesh recovery on the COCO-UniHuman val set. HQNets are jointly trained with all HCP tasks (D+S+P+C). $\downarrow$ means lower is better.
Model Backbone Pose (Mesh.)
[…]
ROMP [68] R-50 119.52 72.27
HQNet w/o GaSM R-50 87.00 54.92
HQNet R-50 84.74 50.80
HQNet ViT-L 76.31 48.26
@@ 5.1 Dataset and Evaluation Metric
COCO-UniHuman Dataset. Our model training exclusively employs COCO-UniHuman train data (in addition to ImageNet pre-training). We follow DINO [85] for augmentation and adopt the 100-epoch training schedule. Model evaluation takes place on COCO-UniHuman val set (2693 images). Due to limitations of the COCO test-dev evaluation server, which lacks support for “Person" category evaluation and attribute recognition, we mainly report results on the val set. The evaluation is based on the standard COCO metrics including Average Precision (AP), $\text{AP}^{M}$ for medium-sized persons and $\text{AP}^{L}$ for large-sized persons. Following [89, 90], we exclude the Small category persons during evaluation due to the lack of annotations in COCO. For attribute recognition, we also use AP with Age-10 metric for evaluation, where the age estimation is considered correct if the prediction error is no larger than 10. For human mesh recovery, we evaluate pose accuracy using MPJPE (Mean Per Joint Position Error) w.r.t. root relative poses and PA-MPJPE (Procrustes-Aligned MPJPE), which is MPJPE calculated after rigid alignment of predicted pose with the ground truth.
OCHuman Dataset [89] is a large benchmark that focuses on heavily occluded humans. It contains no training samples and is intended solely for evaluation purposes. Following [89], we train models on the COCO train set and evaluate models on OCHuman val set (4731 images) and test set (8110 images).
5.2 Results on COCO-UniHuman Dataset
We compare our method to task-specific and multi-task HCP models on the COCO-UniHuman dataset in Table 2. Our models outperform multi-task HCP models and achieves very competitive results against task-specific HCP models. Details about the baselines can be found in Sec. A5. “D”, “S”, “P”, “C” mean model training with Detection (Det.), Segmentation (Seg.), Pose and Classification (Cls.) task respectively.
@@ 5.2 Results on COCO-UniHuman Dataset
Det. Seg. Pose (Kpt.) Det. Seg. Pose (Kpt.)
$\diamondsuit$ Mask R-CNN [20] R-50-FPN - 16.3 - 16.9
$\diamondsuit$ ${}^{\dagger}$ [78] R-50 37.8 30.4
$\diamondsuit$ Pose2Seg [89] R-50-FPN 22.2 28.5 23.8 30.3
$\heartsuit$ ${}^{\dagger}$ [54] Hourglass 32.1 29.5
$\heartsuit$ ${}^{\dagger}$ [26] Hourglass 35.6 34.8
$\heartsuit$ ${}^{\dagger}$ [18] HRNet-w32 37.9 36.5
$\heartsuit$ ${}^{\dagger}$ [10] HRNet-w32 40.0 39.4
$\bigstar$ YOLACT [5] R-101-FPN 13.2 13.5
$\bigstar$ CondInst [72] R-50-FPN 20.3 20.1
$\bigstar$ LSNet-5 [90] DLA-34 25.0 24.9
$\bigstar$ ${}^{\dagger}$ [81] HRNet-w32 39.0 38.1
$\bigstar$ ${}^{\dagger}$ [73] R-50-FPN 29.2 28.3
$\bigstar$ ${}^{\dagger}$ [73] HRNet-w32 44.9 44.0
@@
Comparison with multi-task HCP methods. Pose2Seg [89] is a two-stage human pose-based instance segmentation approach. It uses a standalone keypoint detector for pose estimation and employs human skeleton features for top-down instance segmentation guidance. MultiPoseNet [1] and PersonLab [59] follow bottom-up strategies. CenterNet [94], LSNet [90], and UniHead [40] [^1] UniHead trains separate models for different HCP tasks. are single-stage alternatives. Our R-50 model achieves superior performance in multi-task HCP, without bells and whistles.
Effect of multi-task co-learning. In Table 2, we also compare with different variants of HQNet for various task composition (i.e. D, S, P, C). We observed that co-learning with multiple human-centric tasks leads to improved overall performance. This enhancement can be attributed to the inter-task synergy that arises from jointly training different HCP tasks.
@@ 5.3 Results on the OCHuman Dataset
To verify the performance of HQNet in challenging crowded scenarios, we compare it with recent works on OCHuman dataset [89], which is a crowded scene benchmark for human detection, segmentation, and pose estimation. We show that our model outperforms previous methods under the same ResNet50 backbone network by a large margin. For instance, it outperforms SBL by 9.6 keypoint AP and CondInst by 11.0 segmentation AP on test set. It even achieves superior performance than HrHRNet (40.3 vs 40.0) and LOGO-CAP (40.3 vs 39.0) even with a much smaller backbone (ResNet-50 vs. HRNet-w32). With a stronger backbone, i.e. ViT-L, our HQNet sets new state-of-the-art results on detection (35.8 AP), segmentation (38.8 AP), and pose estimation (45.6 AP).
@@
Table 6: Unseen-task evaluation on PoseTrack21 [14]. ‘FT’ means fine-tuning on PoseTrack21. Our models are evaluated without training on MOT.
@@
Table 6: Unseen-task evaluation on PoseTrack21 [14]. ‘FT’ means fine-tuning on PoseTrack21. Our models are evaluated without training on MOT.
@@ 5.6 More Analysis
Computation cost analysis. In Fig 4, we report the number of parameters of our Res50-based HQNet model variants of different task composition. In HQNet, multiple tasks share the computation cost of the backbone, transformer encoder and decoder. The overhead of each task-specific head is negligible, showing good scalability of HQNet in terms of increasing the number of tasks. It is noteworthy that our model is efficient and its cost is comparable to the task-specific HCP models (e.g. MaskDINO [36]).
Effect of Gender-aided human Model Selection (GaMS) In Table 3, we analyze the effect of Gender-aided human Model Selection (GaMS). We find that incorporating the obtained gender information can assist in selecting proper 3D model of the human body, resulting in more accurate human mesh recovery.
Effect of HQ-Ins Matching. “w/o HQ-Ins Matching” means using detection loss only for bipartite matching [85]. “w/ HQ-Ins Matching” means comprehensively using detection, pose, and segmentation loss for bipartite matching. As shown in Fig. 4, with detection only matching [85], there may be some erroneous cases when one person’s pose is matched to another person. HQ-Ins Matching avoids such errors by comprehensively considering multiple tasks as a whole. More quantitative evaluation can be found in Sec. A3.2.
Figure 3: Computation cost analysis validates the efficiency of HQNet.
Figure 3: Computation cost analysis validates the efficiency of HQNet.
Figure 4: Effect of HumanQuery-Instance (HQ-Ins) Matching.
@@ 6 Conclusion
Limitations and future work. While our work focuses on RGB image based HCP tasks, tasks about video data or multi-modality data (e.g. IR and Depth) also hold significant potential. We encourage future research to explore more comprehensive multi-task human-centric perception.
Acknowledgement. This paper is partially supported by the National Key R&D Program of China No.2022ZD0161000 and the General Research Fund of Hong Kong No.17200622.
@@
Figure A1: Statistics of the COCO-UniHuman benchmark. (a) The gender distribution of COCO-UniHuman is biased towards male. (b) The age distribution ranges from [1, 84] and is biased towards young adults, since images are from public Internet repositories.
In Fig. A1, we show statistics of our proposed COCO-UniHuman dataset. The plots show the distribution of the gender and the apparent age. We find gender and age biases existed in the widely used COCO dataset. The occurrence of men is significantly higher than women in COCO dataset. More specifically, male to female ratio is about 65:35. In addition, the dataset has an unbalanced age distribution. The apparent age distribution ranges from [1, 84], and it is mainly concentrated between the ages of 25 and 35. Analyzing and addressing the gender and age bias in the computer vision system can also be an important topic in the AI community. Future work could also use our benchmark dataset to comprehensively measure and analyze such biases, but it is out of the scope of this paper.
@@
Multi-task co-learning can mitigate over-fitting. From Fig. A2, we observe that training task-specific models on “Person” category only will easily lead to over-fitting problem, the performance decreases with increasing number of epochs. Specifically, in this experiment, we compare the common 1x, 2x, and 4x training settings for RCNN-based methods (i.e. Faster-RCNN, Mask RCNN), and compare 50-epoch and 100-epoch settings for DETR-based methods (i.e. DINO, Mask DINO, and our HQNet). The models are trained using MMDetection [8] with suggested hyper-parameters. We report the Average Precision (AP) for both human detection (solid lines) and the human instance segmentation (dashed lines) on COCO-UniHuman val set. Interestingly, we find that our presented multi-task co-learning (HQNet) can mitigate the over-fitting problem, and the performance consistently improves with the increasing training epochs, demonstrating good scalability.
@@
A3.1 General class models vs person-specific models
@@
Table A4: Effect of HumanQuery-Instance Matching. Experiments are conducted on COCO-UniHuman val set using the ResNet-50 backbone with 100-epoch training setting. We report AP for the “Person” category without Small category person.
Matching Det. Seg. Pose (Kpt.) Cls. (Gender) Cls. (Age)
@@
Figure A3: Qualitative results on COCO-UniHuman val dataset. Our HQNet achieves accurate human detection, human pose estimation, human instance segmentation, human attribute recognition and human mesh estimation simultaneously.
@@
Figure A4: Qualitative results on PoseTrack21 val video dataset. Our HQNet makes a comprehensive human analysis system that can achieve multiple functions: multiple object tracking with human pose estimation, human instance segmentation, and human attribute recognition. Our HQNet is only trained on the COCO-UniHuman image-based dataset without finetuning on the PoseTrack21 video-based dataset.
@@
In Fig. A3 and Fig. A4, we show some qualitative results of HQNet with ResNet-50 backbone. In Fig. A3, we show some qualitative results on COCO-UniHuman val dataset for human detection, human pose estimation, human instance segmentation and human attribute recognition, and human mesh estimation. The model can recognize the gender and age of different people In Fig. A4, we visualize the results of human detection and tracking (same color for same id), human pose estimation, human instance segmentation, gender estimation, age estimation and mesh estimation. As introduced in the section of “Unseen-task generalization” in the main paper, we directly apply our HQNet on multiple object tracking on the challenging PoseTrack21 [14] dataset, where our models are trained only on the COCO-UniHuman image-based dataset without explicitly tuned for multi-object tracking (MOT) on video-based dataset like PoseTrack21. Our learned Human Query, which encodes both spatial and visual cues, can serve as good embedding features to distinguish different human instances. Therefore, our human tracking is robust to heavy occlusion, and the id can recover from occlusions. Our HQNet makes a comprehensive all-in-one human analysis system that can achieve multiple functions: multiple object tracking with human pose estimation, human instance segmentation, human attribute recognition and human mesh estimation.
Figure A5: Visualization of deformable attention sampling points. The results are from models trained for different HCP tasks. Left: detection and segmentation. Middle: detection and pose. Right: detection, pose, segmentation and attribute.
@@
In Fig. A5, we visualize the sampling locations of deformable attention for different HCP models. We show the results of the last decoder layer in HQNet-ResNet50. Each sampling point is marked as a red-filled circle. The left results are from the model trained for detection and segmentation ($M_{D+S}$). The middle ones are from the model trained for detection and pose ($M_{D+P}$). And the right ones are from the model trained for detection, segmentation, pose and attribute ($M_{D+P+S+C}$). With the segmentation task, we notice that some of the sampling points of $M_{D+S}$ are distributed near the boundary of the human body, and some are distributed in the background to capture more context information. The sampling points of $M_{D+P}$ have higher probability to distribute inside the human body and some of the points are located closer to the defined human body keypoints, especially the face, arms, and legs. $M_{D+P+S+C}$ combines the characteristics of $M_{D+S}$ and $M_{D+P}$.
@@
In this work, we jointly train multiple human-centric perception (HCP) tasks, including human detection, human instance segmentation, human pose estimation human attribute (gender and age) recognition and human mesh estimation.
For human detection, we follow DINO [85] to apply focal loss [44] for classification $L^{focal}_{cls}$ and detection loss (L1 regression loss $L^{reg}_{det}$ and GIOU loss [61] $L^{giou}_{det}$). For human pose estimation, we follow PETR [65] to use focal loss for classifying valid and invalid human instances $L^{focal}_{kpt}$, L1 keypoint regression loss $L^{reg}_{kpt}$, OKS loss $L^{oks}_{kpt}$, and auxiliary heatmap loss $L^{hm}_{kpt}$. For segmentation loss, we use binary cross-entropy loss $L^{bce}_{seg}$ and dice loss $L^{dice}_{seg}$. For attribute recognition, we have binary cross-entropy loss for gender estimation $L^{bce}_{gender}$ and mean-variance loss [58] for age estimation $L^{mean}_{age}$ and $L^{var}_{age}$. For mesh estimation, we use L1 regression loss for pose estimation $L^{reg}_{pose}$, shape estimation $L^{reg}_{shape}$ and the 3D joints regressed from the body model $L^{reg}_{3d}$.
@@
$+\lambda^{focal}_{2d}L^{focal}_{2d}+\lambda^{reg}_{2d}L^{reg}_{2d}+\lambda^{oks}_{2d}L^{oks}_{2d}+\lambda^{hm}_{2d}L^{hm}_{2d}$
$+\lambda^{reg}_{pose}L^{reg}_{pose}+\lambda^{reg}_{shape}L^{reg}_{shape}+\lambda^{reg}_{3d}L^{reg}_{3d}$
@@
$+\lambda^{bce}_{gender}L^{bce}_{gender}+\lambda^{mean}_{age}L^{mean}_{age}+\lambda^{var}_{age}L^{var}_{age},,$
@@
Pose $\lambda^{focal}_{kpt}$ 1.0
@@
We follow the setting of DINO [85] to augment the input image by random crop, random flip, and random resize. Specifically, we randomly resize the input image to have its shorter side between 480 and 800 pixels and its longer side less or equal to 1333. The models are trained with AdamW optimizer [31] with base learning rate of $1\times 10^{-4}$, momentum of $0.9$ and weight decay of $1\times 10^{-4}$. For all experiments, the models are trained for 100 epochs with a total batch size of 16 and the initial learning rate is decayed at 80th epoch by a factor of 0.1. We use 16 Tesla V100 GPUs for model training.
@@
Details about human instance segmentation baselines. For human instance segmentation, we contrast HQNet with state-of-the-art general and human-specific instance segmentation methods. Mask R-CNN [20] is an end-to-end top-down approach that optimizes object detection and instance segmentation jointly. Given our one-stage pipeline, we also compare against one-stage methods , including PolarMask [79], MEInst [87], YOLACT [5], and CondInst [72]. Results of PolarMask, MEInst, YOLACT, CondInst are from [90], which are obtained by re-training and evaluating the models on COCO “Person” category only. PolarMask encodes the instance mask with coordinates, while MEInst encodes the mask into a compact representation vector. YOLACT and CondInst use a series of global prototypes and linear coefficients to represent instance masks. Instead we learn instance-aware Human Query to decouple each human instance.
@@
Details about mesh estimation baseline. For human mesh estimation, we compare HQNet with state-of-the-art one-stage monocular method ROMP [68]. Following official settings, we train ROMP on COCO-UniHuman with ResNet50 backbone.